OpenStudy (anonymous):

k/11 +5/13 = -10/13. Please help I been trying to solve this for an hour

OpenStudy (anonymous):

k/11+5/13=-10/13 -5/13 -5/13 k/11=-15/13 *11 * 11 k=-12 9/13

hero (hero):

\(\frac{k}{11} + \frac{5}{13} = -\frac{10}{13}\) Add 10/13 to both sides to get \(\frac{k}{11} + \frac{5}{13} + \frac{10}{13} = 0\) Combine the second and third fractions to get: \(\frac{k}{11} + \frac{15}{13}= 0\) Subtract 15/13 from both sides to get: \(\frac{k}{11} = - \frac{15}{13}\) Cross mutliply to get 13k = -11(15) 13k = 165 Divide both sides by 13 \(k = \frac{165}{13}\)

OpenStudy (anonymous):

eh his is probably right.

hero (hero):

No, mine is wrong because I forgot the negative.

hero (hero):

It's right if you include the negative

OpenStudy (anonymous):

oh, ok cool.

hero (hero):

\(k = -\frac{165}{13}\)

OpenStudy (anonymous):

Thanks